A public project consisting of a construction public ice rink is being considered. project will require initial construction cost of 1.5 million and subsequest maint with estimated annual cost of 69000. Project is assumed to last 19 years. After project id over , teh facility and land will be sold to private company.for an estimated 301000 The annual benefits to public are estimated 264000 per year. Assuming 7% cost of capital, calculate modified B-C ratio of project. Be careful to use correct formulae for computing b-c ratio
Answer #1

Benefit – cost ratio = present value of cash benefits / present value of cost

Annual cash inflow = annual benefits – annual cost = 264000-69000 = 195000

PV factor

Annual cash inflow

195000

10.33560

2015442

Salvage value

301000

0.27651

83230

Total present value of cash inflows

$2098672

PV of ordinary annuity of $1 @ n=19, i=7% is 10.33560

PV of $1 @ n=19, i=7% is 0.27651

Benefit – cost ratio = 2098672/1500000 = 1.40